calculate the density, specific weight and weight of 1.5 liters of gasoline with a specific gravity of 0.70

1
Expert's answer
2022-02-24T11:30:46-0500

Answer

Sg=0.70

Volume

V=1.5"*10^{-3}m^3"

Let mass is m kg

So density

"d=Sg.\rho=0.7*1000\\=700kg/m^3"

Specific weight

"=dg\\=700*9.81\\=6867N/m^3"